# Revisiting Stochastic Proximal Point Methods: Generalized Smoothness and Similarity

### v1

@@ 7 Experiments
In the second set of experiments, we investigate the dependence of convergence on the initial point. The objective is to verify that the convergence of SPPM-inexact is independent of the initial point and to assess the tightness of the theoretical analysis, i.e., whether the convergence rate depends on the distance between the initial and optimal points. For this experiment, we analyze three different values of $s$ ($s=2,3,4$) and, for each case, randomly select five different initial points with norms $\|x_{0}\|=0.1,1,10,50,100$, while keeping all other parameters unchanged. As shown in Figure 2, the convergence rates are nearly identical across all cases, suggesting that the upper bound $\|x_{k}-x_{k+1}\|^{2}\leq\|x_{0}-x_{*}\|^{2}$ provided in Lemma 5.1 may be overly restrictive for certain problems. Further investigation of this effect is a promising direction for future research.
@@
The final set of experiments investigates the effect of varying the maximum number of iterations for the inner solver, without enforcing a stopping condition based on the gradient norm. The purpose of this analysis is to demonstrate that if the inner solver fails to achieve sufficient accuracy in solving the proximal step, the method may either diverge or exhibit slower convergence. We conduct two types of experiments. In the first, we consider functions with different values of $s$, ranging from $2$ to $20$, and run the algorithm with varying maximum iteration limits for the inner solver: $T=5,20,50,100,500$. The number of functions is set to $n=100$, and the dimension is $d=5$. For the second experiment in this set, we modify the original problem to ensure strong convexity of $f$, enabling us to verify that the convergence rate becomes linear in this setting. The modified problem (2) is defined as follows: $f_{i}(x)=a_{i}\|x\|^{2s}+\lambda\langle e_{i},x\rangle^{2},$ where $a_{i}\in\mathbb{R}^{+}$ for all $i=1,\dots,n$, $\lambda\in\mathbb{R}^{+}$, and $e_{i}$ is a unit vector with its $i$-th coordinate equal to one. In this case, each function $f_{i}$ is $(L_{0},L_{1})$-smooth, with $L_{0}=2s+2\lambda$ and $L_{1}=2s-1$, and convex. Furthermore, $f$ is $\frac{\lambda}{n}$-strongly convex. We set $n=d=100$, $\lambda=2$, and $s=2$, and conduct experiments with inner solver iteration limits of $1,10,15,100,200$.
@@ 8 Conclusion
Acknowledgments
This work was supported by funding from King Abdullah University of Science and Technology (KAUST): i) KAUST Baseline Research Scheme, ii) Center of Excellence for Generative AI, under award number 5940, iii) SDAIA-KAUST Center of Excellence in Artificial Intelligence and Data Science.
@@ Appendix A Fundamental Lemmas
[…]
@@
[…]
@@
where $\hat{x}^{k}$ is a vector chosen from the set of iterates $x_{0}$, . . . , $x_{k-1}$ uniformly at random.
@@
Using the identity $-2\langle a,b\rangle=-\|a+b\|^{2}+\|a\|^{2}+\|b\|^{2}$, we rewrite the expression as
@@
$f_{\xi_{k}}(x_{*})\geq f_{\xi_{k}}(\hat{x}_{k})+\langle\nabla f_{\xi_{k}}(\hat{x}_{k}),x_{*}-\hat{x}_{k}\rangle.$
which implies:
@@
$\|x_{k+1}-x_{*}\|^{2}$ $\leq\|x_{k}-x_{*}\|^{2}-2\gamma(f_{\xi_{k}}(x_{k}^{\Psi})-f_{\xi_{k}}(x_{*}))-\|x_{k}^{\Psi}-x_{k}\|^{2}+\|\hat{x}_{k}-x_{k}\|^{2}-\|x_{k}-\hat{x}_{k}\|^{2}+\gamma^{2}\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}$
$=\|x_{k}-x_{*}\|^{2}-2\gamma(f_{\xi_{k}}(x_{k}^{\Psi})-f_{\xi_{k}}(x_{*}))-\|x_{k}^{\Psi}-x_{k}\|^{2}+\gamma^{2}\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}.$ (21)
@@
Appendix C Proof of Theorems 6.1 and 6.2
@@
Proof.
Define
$\bar{x}_{k+1}=\mathbb{E}[x_{k+1}|\mathcal{F}_{k}].$ (26)
Then
[…]
$\quad-\mathbb{E}\left[\|x_{k}-x_{k+1}\|^{2}|\mathcal{F}_{k}\right]$
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $=\|x_{k}-x_{*}\|^{2}-2\gamma\left(f(\bar{x}_{k+1})-f(x_{*})+\mathbb{E}\left[\langle\nabla f_{\xi_{k}}(\bar{x}_{k+1})-\nabla f(\bar{x}_{k+1}),x_{k+1}-\bar{x}_{k+1}\rangle|x_{k}\right]\right)$
$\quad-\mathbb{E}\Big{[}\|\gamma(\nabla f_{\xi_{k}}(\bar{x}_{k+1})-\nabla f(\bar{x}_{k+1}))+(x_{k+1}-\bar{x}_{k+1})\|^{2}-$
$\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}-\|x_{k}-\bar{x}_{k+1}\|^{2}+\gamma^{2}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(\bar{x}_{k+1})-\nabla f(\bar{x}_{k+1})\|^{2}|\mathcal{F}_{k}\right].$
Finally, using the star similarity condition, we obtain
Taking the full expectation, we obtain
Applying this inequality recursively, we get
For convenience, we restate Theorem 6.2 here:
Theorem C.2.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.4 (Interpolation) and 2.8 (Star Similarity) hold. Consider SPPM-inexact with Assumption 2.9 satsified. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c<1$, then the iterates of SPPM-inexact satisfy, for every $k\geq 0$,
$\mathbb{E}\left[\|x_{k}-x_{*}\|^{2}\right]\leq\left(1\!-\!\min\left(\frac{\gamma\mu}{4},\frac{1\!-\!c}{2}\right)\right)^{k}\|x_{0}-x_{*}\|^{2}.$
Proof.
To avoid repetitions, we start from (22):
Instead of (26), we define $\bar{x}_{k+1}=\mathbb{E}[x^{\Psi}_{k}|x_{k}]$.
Using convexity of $f_{\xi_{k}}$, we get
$\|x_{k+1}-x_{*}\|\leq\|x_{k}-x_{*}\|^{2}-2\gamma(f_{\xi_{k}}(\bar{x}_{k+1})+\langle\nabla f_{\xi_{k}}(\bar{x}_{k+1}),x^{\Psi}_{k}-\bar{x}_{k+1}\rangle-f_{\xi_{k}}(x_{*}))-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})\|x_{k}-x^{\Psi}_{k}\|^{2}.$
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-2\gamma\mathbb{E}\left[f_{\xi_{k}}(\bar{x}_{k+1})+\langle\nabla f_{\xi_{k}}(\bar{x}_{k+1}),x^{\Psi}_{k}-\bar{x}_{k+1}\rangle-f_{\xi_{k}}(x_{*})|x_{k}\right]$
$\quad-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})E\left[\|x_{k}-x^{\Psi}_{k}\|^{2}|x_{k}\right]$
Since $\mathbb{E}\left[\langle f(\bar{x}_{k+1}),x_{k+1}-\bar{x}_{k+1}\rangle|x_{k}\right]=0$, we can write
Let $a\in(0,1)$. Then
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-2\gamma\left(f(\bar{x}_{k+1})-f(x_{*})\right)-2\frac{1}{a}\mathbb{E}\left[\langle\gamma(\nabla f_{\xi_{k}}(\bar{x}_{k+1})-f(\bar{x}_{k+1})),a(x^{\Psi}_{k}-\bar{x}_{k+1})\rangle|x_{k}\right]$
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})\mathbb{E}\left[\|x^{\Psi}_{k}-\bar{x}_{k+1}\|^{2}|x_{k}\right]-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})\|x_{k}-\bar{x}_{k+1}\|^{2}$
$\quad-\frac{1}{a}\mathbb{E}\left[\|\gamma(\nabla f_{\xi_{k}}(\bar{x}_{k+1})-\nabla f(\bar{x}_{k+1}))+(x_{k+1}-\bar{x}_{k+1})\|^{2}|x_{k}\right]$
By setting $a=(1-\frac{\eta\gamma^{2}}{T^{\alpha}})$, we obtain
Now, using the star similarity assumption (2.8), we get
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})\|x_{k}-\bar{x}_{k+1}\|^{2}+\frac{\gamma^{2}\delta_{*}^{2}}{(1-\frac{\eta\gamma^{2}}{T^{\alpha}})}\|\bar{x}_{k+1}-x_{*}\|^{2}.$
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}|x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}-(1-c)\|x_{k}-\bar{x}_{k+1}\|^{2}+2\gamma^{2}\delta_{*}^{2}\|\bar{x}_{k+1}-x_{*}\|^{2}.$
We want $2\gamma^{2}\delta_{*}^{2}\leq\frac{\gamma\mu}{2}$, which means $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$, so we get
$\leq\|x_{k}-x_{*}\|^{2}-\min\left(\frac{\gamma\mu}{4},\frac{1-c}{2}\right)\|x_{k}-x_{*}\|^{2}$
$=\left(1-\min\left(\frac{\gamma\mu}{4},\frac{1-c}{2}\right)\right)\|x_{k}-x_{*}\|^{2}.$
Taking the full expectation, we obtain
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\right]\leq\left(1-\min\left(\frac{\gamma\mu}{4},\frac{1-c}{2}\right)\right)\mathbb{E}\left[\|x_{k}-x_{*}\|^{2}\right].$
Iterating over $k$, the result follows:
$\mathbb{E}\big{[}\|x_{k+1}-x_{*}\|^{2}\big{]}\leq\left(1-\min\left(\frac{\gamma\mu}{4},\frac{1-c}{2}\right)\right)^{k+1}\|x_{0}-x_{*}\|^{2}.$
Appendix D Proof of Theorems 6.4 and 6.5
For convenience, we restate Theorem 6.4 here:
Theorem D.1.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.3(Strong convexity of $f$), 2.8 (Star Similarity) and 6.3 (Bounded Variance at Optimum) hold. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$, we have, for every $k\geq 0$,
@@
Since $\mathbb{E}\left[\langle f(\bar{x}_{k+1}),x_{k+1}-\bar{x}_{k+1}\rangle|\mathcal{F}_{k}\right]=0$, we can write:
@@
By substituting $s=1-a$ and choosing $a=\frac{1}{2}$, we get
@@
Theorem D.2.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.8 (Star Similarity) and 6.3 (Bounded Variance) hold. Consider SPPM-inexact with Assumption 2.9 satisfied. If the stepsize satisfies $\gamma\leq\frac{\mu(1-c)}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c$, where $10<c<$ 1 is a constant, then SPPM-inexact satisfies, for every $k\geq 0$,
$\mathbb{E}\big{[}\|x_{k}-x_{*}\|^{2}\big{]}\leq\left(1-\frac{1}{4}\min\left(\gamma\mu,1\right)\right)^{k}\|x_{0}-x_{*}\|^{2}+\max\left(\frac{1}{\gamma\mu},1\right)\frac{8\gamma^{2}\sigma_{*}^{2}}{(1-c)}.$
@@
$\quad+\frac{\gamma^{2}}{s}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]+sE\left[\|\bar{x}_{k+1}-x^{\Psi}_{k}\|^{2}\mid x_{k}\right]-(1-\frac{\eta\gamma^{2}}{T^{\alpha}})\mathbb{E}\left[\|x^{\Psi}_{k}-\bar{x}_{k+1}\|^{2}\mid x_{k}\right]$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\mid x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}+\frac{2\gamma^{2}\delta_{*}^{2}}{1-\frac{\eta\gamma^{2}}{T^{\alpha}}}\|\bar{x}_{k+1}-x_{*}\|^{2}+\frac{2\gamma^{2}}{1-\frac{\eta\gamma^{2}}{T^{\alpha}}}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\mid x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\frac{\gamma\mu}{2}\|\bar{x}_{k+1}-x_{*}\|^{2}-\frac{1}{2}\|x_{k}-\bar{x}_{k+1}\|^{2}+\frac{2\gamma^{2}}{1-c}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]$
$\leq\|x_{k}-x_{*}\|^{2}-\min\left(\frac{\gamma\mu}{4},\frac{1}{4}\right)\|x_{k}-\bar{x}_{*}\|^{2}+\frac{2\gamma^{2}}{1-c}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\right]\leq\left(1-\min\left(\frac{\gamma\mu}{4},\frac{1}{4}\right)\right)\mathbb{E}[\|x_{k}-x_{*}\|^{2}]+\frac{2\gamma^{2}\sigma_{*}^{2}}{1-c}.$

### v2

@@ 7.2.2 Dependence on the Initial Point
Finally, we study the dependence of convergence on the initial point. The objective is to verify that the convergence of SPPM-inexact is independent of the initial point and to assess the tightness of the theoretical analysis, i.e., whether the convergence rate depends on the distance between the initial and optimal points. For this experiment, we analyze three different values of $s$ ($s=2,3,4$) and, for each case, randomly select five different initial points with norms $\|x_{0}\|=0.1,1,10,50,100$, while keeping all other parameters unchanged. As shown in Figure 4, the convergence rates are nearly identical across all cases, suggesting that the upper bound $\|x_{k}-x_{k+1}\|^{2}\leq\|x_{0}-x_{*}\|^{2}$ provided in Lemma 5.1 may be overly restrictive for certain problems. Further investigation of this effect is a promising direction for future research.
@@ 8 Conclusion
Acknowledgements
This work was supported by funding from King Abdullah University of Science and Technology (KAUST):
i) KAUST Baseline Research Scheme,
ii) Center of Excellence for Generative AI (award no. 5940),
iii) Competitive Research Grant (CRG) Program (award no. 6460),
iv) SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence (SDAIA-KAUST AI).
@@ Appendix A Fundamental Lemmas
$f_{\xi}(x)-f_{\xi}(y)$ $=\int_{0}^{1}\langle\nabla f_{\xi}(y+t(x-y)),x-y\rangle\,dt$
$=\int_{0}^{1}\langle\nabla f_{\xi}(y+t(x-y))-\nabla f_{\xi}(y),x-y\rangle\,dt+\langle\nabla f_{\xi}(y),x-y\rangle.$
@@
$D_{f_{\xi}}(x,y)$ $=\int_{0}^{1}\langle\nabla f_{\xi}(y+t(x-y))-\nabla f_{\xi}(y),x-y\rangle\,dt$
$\leq\int_{0}^{1}\|\nabla f_{\xi}(y+t(x-y))-\nabla f_{\xi}(y)\|\,\|x-y\|\,dt.$
@@
From Proposition 3.2 (Point 2) in Chen et al. (2023), the function $f_{\xi}\colon\mathbb{R}^{d}\to\mathbb{R}$ is symmetrically $(L_{0},L_{1})$-smooth if and only if for all $x,y\in\mathbb{R}^{d}$,
$\|\nabla f_{\xi}(x)-\nabla f_{\xi}(y)\|\leq\left(L_{0}+L_{1}\|\nabla f_{\xi}(y)\|\right)\exp\left(L_{1}\|x-y\|\right)\|x-y\|.$
Hence, defining
$\phi\left(\|x-y\|,\|\nabla f_{\xi}(y)\|\right):=\left(L_{0}+L_{1}\|\nabla f_{\xi}(y)\|\right)\exp\left(L_{1}\|x-y\|\right),$
we obtain
$\|\nabla f_{\xi}(x)-\nabla f_{\xi}(y)\|\leq\phi\left(\|x-y\|,\|\nabla f_{\xi}(y)\|\right)\|x-y\|,$
which confirms that $f_{\xi}$ is $\phi$-smooth according to Assumption 4.1.
A.3 Proof of Lemma 4.4
From Proposition 3.2 (Point 1) in Chen et al. (2023), the function $f_{\xi}\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to satisfy the $\alpha$-symmetric generalized-smoothness condition if and only if for all $x,y\in\mathbb{R}^{d}$,
$\|\nabla f_{\xi}(x)-\nabla f_{\xi}(y)\|\leq\|x-y\|\left(K_{0}+K_{1}\|\nabla f_{\xi}(y)\|^{\alpha}+K_{2}\|x-y\|^{\frac{\alpha}{1-\alpha}}\right),$
where the constants $K_{0},K_{1},K_{2}$ are defined as:
$K_{0}$ $:=L_{0}\left(2^{\frac{\alpha^{2}}{1-\alpha}}+1\right),$
$K_{1}$ $:=L_{1}\cdot 2^{\frac{\alpha^{2}}{1-\alpha}}\cdot 3^{\alpha},$
$K_{2}$ $:=L_{1}^{\frac{1}{1-\alpha}}\cdot 2^{\frac{\alpha^{2}}{1-\alpha}}\cdot 3^{\alpha}(1-\alpha)^{\frac{\alpha}{1-\alpha}}.$
Hence, defining
$\phi\left(\|x-y\|,\|\nabla f_{\xi}(y)\|\right):=K_{0}+K_{1}\|\nabla f_{\xi}(y)\|^{\alpha}+K_{2}\|x-y\|^{\frac{\alpha}{1-\alpha}},$
we obtain
$\|\nabla f_{\xi}(x)-\nabla f_{\xi}(y)\|\leq\phi\left(\|x-y\|,\|\nabla f_{\xi}(y)\|\right)\|x-y\|,$
which confirms that $f_{\xi}$ is $\phi$-smooth in the sense of Assumption 4.1.
A.4 Proof of Lemma 5.1
@@
where $\hat{x}_{k}$ is a vector chosen from the set of iterates $x_{0}$, . . . , $x_{k-1}$ uniformly at random.
@@
Using the identity $2\langle a,b\rangle=-\|a+b\|^{2}+\|a\|^{2}+\|b\|^{2}$, we rewrite the expression as
@@
$f_{\xi_{k}}(x_{*})\geq f_{\xi_{k}}(\hat{x}_{k})+\langle\nabla f_{\xi_{k}}(\hat{x}_{k}),x_{*}-\hat{x}_{k}\rangle,$
which implies
@@
$\|x_{k+1}-x_{*}\|^{2}$ $\leq\|x_{k}-x_{*}\|^{2}-2\gamma(f_{\xi_{k}}(x_{k}^{\Psi})-f_{\xi_{k}}(x_{*}))-\|x_{k}^{\Psi}-x_{k}\|^{2}+\|\hat{x}_{k}-x_{k}\|^{2}-\|x_{k}-\hat{x}_{k}\|^{2}$
$\quad+\gamma^{2}\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}$
$=\|x_{k}-x_{*}\|^{2}-2\gamma(f_{\xi_{k}}(x_{k}^{\Psi})-f_{\xi_{k}}(x_{*}))-\|x_{k}^{\Psi}-x_{k}\|^{2}+\gamma^{2}\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}.$
@@
B.2.1 Practical details (parameter-free Armijo on $\Psi_{k}$).
Fix $k$. Recall $\Psi_{k}$ and $x_{k}^{\Psi}$ from (4). By the optimality condition,
$x_{k}^{\Psi}+\gamma\nabla f_{\xi_{k}}(x_{k}^{\Psi})=x_{k}\;\;\Longrightarrow\;\;\|\nabla f_{\xi_{k}}(x_{k}^{\Psi})\|=\frac{1}{\gamma}\|x_{k}-x_{k}^{\Psi}\|.$ (24)
From the proof of Theorem 5.4,
$\|x_{k}-x_{k}^{\Psi}\|\leq\|x_{0}-x_{*}\|=:R,$ (25)
hence by (24),
$\|\nabla f_{\xi_{k}}(x_{k}^{\Psi})\|\leq\frac{R}{\gamma}.$ (26)
Let $B_{k}:=\{z:\|z-x_{k}^{\Psi}\|\leq R\}$. For $y\in B_{k}$, by $\phi$-smoothness and (26),
$\|\nabla f_{\xi_{k}}(y)-\nabla f_{\xi_{k}}(x_{k}^{\Psi})\|$ $\leq\phi(\|y-x_{k}^{\Psi}\|,\|\nabla f_{\xi_{k}}(x_{k}^{\Psi})\|)\|y-x_{k}^{\Psi}\|$
$\leq\phi(R,\tfrac{R}{\gamma})R.$ (27)
Thus
$\|\nabla f_{\xi_{k}}(y)\|\leq G:=\frac{R}{\gamma}+\phi(R,\tfrac{R}{\gamma})R,\qquad y\in B_{k}.$ (28)
Now for any $x,y\in B_{k}$,
$\|\nabla f_{\xi_{k}}(x)-\nabla f_{\xi_{k}}(y)\|$ $\leq\phi(\|x-y\|,\|\nabla f_{\xi_{k}}(y)\|)\|x-y\|$
$\leq\phi(2R,G)\|x-y\|.$ (29)
Since $\nabla\Psi_{k}(x)=\nabla f_{\xi_{k}}(x)+\tfrac{1}{\gamma}(x-x_{k})$, we obtain
$\|\nabla\Psi_{k}(x)-\nabla\Psi_{k}(y)\|\leq(\phi(2R,G)+\tfrac{1}{\gamma})\|x-y\|=:L_{\rm loc}\|x-y\|,\quad\forall x,y\in B_{k}.$ (30)
Here $L_{\rm loc}$ is independent of $k$, and $\Psi_{k}$ is $\mu$–strongly convex with $\mu=\tfrac{1}{\gamma}$.
Implication for Assumption 2.9. Since $\Psi_{k}$ is $\mu$-strongly convex and has a locally Lipschitz gradient on $B_{k}$, gradient descent on $\Psi_{k}$ starting at $x_{k}$ with Armijo backtracking enjoys linear convergence. Hence there exist $\rho\in(0,1)$ and $C>0$ (independent of $k$) such that
$\|x^{t}-x_{k}^{\Psi}\|\leq\rho^{t}\|x_{k}-x_{k}^{\Psi}\|,\qquad\|\nabla\Psi_{k}(x^{t})\|\leq C\,\rho^{t}\|x_{k}-x_{k}^{\Psi}\|.$
Therefore, for any $\alpha>0$,
$\|\nabla\Psi_{k}(x^{T})\|^{2}\leq\eta\,\frac{\|x_{k}-x_{k}^{\Psi}\|^{2}}{T^{\alpha}},$
with $\eta=C^{2}C_{\alpha}$ where $C_{\alpha}$ satisfies $\rho^{2T}\leq C_{\alpha}/T^{\alpha}$ for all $T\geq 1$.
Appendix C Proof of Theorems 6.3 and 6.4
@@
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.3(Strong convexity of $f$), 2.8 (Star Similarity) and 6.1 (Bounded Variance at Optimum) hold. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$, we have, for every $k\geq 0$,
@@
By substituting s=1as=1-a and choosing $a=\frac{1}{2}$, we get
@@
Theorem C.2.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.8 (Star Similarity) and 6.1 (Bounded Variance) hold. Consider SPPM-inexact with Assumption 2.9 satisfied. If the stepsize satisfies $\gamma\leq\frac{\mu(1-c)}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c$, where $0<c<1$ is a constant, then SPPM-inexact satisfies, for every $k\geq 0$,
$\mathbb{E}\big[\|x_{k}-x_{*}\|^{2}\big]\leq\left(1-\frac{1}{2}\min\left(\frac{\gamma\mu}{2},1-c\right)\right)^{k}\|x_{0}-x_{*}\|^{2}+\max\left(\frac{2}{\gamma\mu},\frac{1}{1-c}\right)\frac{4\gamma^{2}\sigma_{*}^{2}}{(1-c)}.$
@@
$\quad+\frac{\gamma^{2}}{s}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]+sE\left[\|\bar{x}_{k+1}-x^{\Psi}_{k}\|^{2}\mid x_{k}\right]$
$\quad-\left(1-\frac{\eta\gamma^{2}}{T^{\alpha}}\right)\mathbb{E}\left[\|x^{\Psi}_{k}-\bar{x}_{k+1}\|^{2}\mid x_{k}\right]-\left(1-\frac{\eta\gamma^{2}}{T^{\alpha}}\right)\|x_{k}-\bar{x}_{k+1}\|^{2}.$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\mid x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\gamma\mu\|\bar{x}_{k+1}-x_{*}\|^{2}+\frac{2\gamma^{2}\delta_{*}^{2}}{1-\frac{\eta\gamma^{2}}{T^{\alpha}}}\|\bar{x}_{k+1}-x_{*}\|^{2}$
$\quad+\frac{2\gamma^{2}}{1-\frac{\eta\gamma^{2}}{T^{\alpha}}}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]-\left(1-\frac{\eta\gamma^{2}}{T^{\alpha}}\right)\|x_{k}-\bar{x}_{k+1}\|^{2}.$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\mid x_{k}\right]$ $\leq\|x_{k}-x_{*}\|^{2}-\frac{\gamma\mu}{2}\|\bar{x}_{k+1}-x_{*}\|^{2}-(1-c)\|x_{k}-\bar{x}_{k+1}\|^{2}+\frac{2\gamma^{2}}{1-c}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]$
$\leq\|x_{k}-x_{*}\|^{2}-\frac{1}{2}\min\left(\frac{\gamma\mu}{2},1-c\right)\|x_{k}-\bar{x}_{*}\|^{2}+\frac{2\gamma^{2}}{1-c}\mathbb{E}\left[\|\nabla f_{\xi_{k}}(x_{*})\|^{2}\mid x_{k}\right]$
@@
$\mathbb{E}\left[\|x_{k+1}-x_{*}\|^{2}\right]\leq\left(1-\frac{1}{2}\min\left(\frac{\gamma\mu}{2},1-c\right)\right)\mathbb{E}[\|x_{k}-x_{*}\|^{2}]+\frac{2\gamma^{2}\sigma_{*}^{2}}{1-c}.$